What is the horizontal displacement of the object thrown into the air at a velocity of 30 m/s with an angle of 28 degree?


1
Expert's answer
2021-12-06T09:50:05-0500

Given:

"v_0=30\:\rm m/s"

"\theta=28^\circ"


The range of projectile

"R=\frac{v_0^2\sin2\theta}{g}=\frac{30^2\sin56^\circ}{9.8}=76\: \rm m"